OpenStudy (donnie1999):

What is the inverse of the function below? f (x) = x - 7 ----- 3

OpenStudy (donnie1999):

@dcskaf

OpenStudy (anonymous):

Same thing replace f(x) with y y=x-7 --- 3 then switch them x=y-7 ---- 3

OpenStudy (anonymous):

First multiply then add the 7

OpenStudy (donnie1999):

so would the answer be f-^1(x)=3x+7 ???

OpenStudy (anonymous):

yep you got it

OpenStudy (donnie1999):

r u sure???

OpenStudy (anonymous):

yeah